# Foregrounding Artist Opinions: A Survey Study on Transparency, Ownership, and Fairness in AI Generative Art

### v4

@@ Abstract.
An illustration of a rose garden connecting the minds of artists and being pruned by a machine
1. Introduction
Art has a rich history in co-evolution with the development of new technologies (Etro and Galenson, 2023). The advent of Artificial Intelligence (AI) tools used to create or assist in the creative process carries great potential utility for artists but also harbors the significant potential for harm to the art workforce and artistic and intellectual property. To examine this tension between utility and threat, we surveyed 459 artists about their views on AI-generated art. The survey results reveal the nuanced perspectives of artists on Generative AI. When the discussion is thorough and directly related to their art practice, it prevents their views from being reduced to a simplistic binary sentiment, which fails to represent the diversity of their field. By capturing artists’ views through a survey, this method moves beyond anecdotal evidence and provides a counterbalance to third-party opinions as well as those of the Generative AI industry.
@@ 1. Introduction
2. Background
There is ongoing disagreement regarding whether the output of Generative AI can be considered art (US Federal Trade Commission, 2023; Jiang et al., 2023); we will not examine the topic in depth here. For convenience, we’ll refer to AI-generated output as “art” in this paper without asserting or refuting its artistic merit.
Jiang et al. define art as a human endeavor linked to culture and experience (Jiang et al., 2023). They argue that human creativity, distinguished by unique inspiration and corporeal engagement, differs from AI-generated work. Recent court decisions in the US have similarly denied copyright protection to autonomously produced AI creations, underscoring the role of human creativity in copyright law (US District Court for the District of Columbia, 2023).
Three compelling arguments against considering AI output as art emerge: art’s human origin devoid of ”organic pressure,” the intimate connection between artist and world, and art’s unique role in human experience (Jiang et al., 2023; Huyghe, 1970; Pfeifer and Bongard, 2006). Regardless of the artistry debate, AI-generated creations can meet commercial standards, like book covers, entering diverse artistic spheres, and impacting practitioners.
Artistic use of technology and AI tools (Cetinic and She, 2022) is not novel; artists historically adopt tools, including machines and automation, often innovating beyond standard applications (Huyghe, 1970). Artists engaging with technology question its role in art and explore its boundaries, raising queries about intentionality, distraction, and artistic agency (Karson, 2023).
It is important to note that, although how technology should be incorporated into art may be up for debate, that technology is part of the artistic process is not new or ambiguous (Mazzone and Elgammal, 2019). Artists have always used tools (Huyghe, 1970), and many art forms involve machines (such as woodwork, metalwork, and digital art) and automation (such as digital printers and cameras). Some artists also choose to use technology in ways other than its intended purpose or develop new art-making techniques by breaking technology. Such artists may be both suspicious and intrigued by the technology they use, questioning the relationship between humans and machines and expanding art practices and processes. When presented with new technology, an artist may ask questions like “How can I keep my work with the machine intentional? How might this technology distract me? Is the machine using me, or am I using it? Is this image-making or art-making? How does this automation serve my artistic intentions and practice?” (Karson, 2023).
@@ 2.1. Generative AI Models
Generative AI models can generate outputs across various media types based on user inputs, such as text prompts or multimedia (Gozalo-Brizuela and Garrido-Merchán, 2023). Key players in textual Generative AI include large language models (LLMs) like OpenAI’s ChatGPT and GPT-4, while visual models like StableDiffusion and Midjourney modify images using iterative noise adjustments guided by diffusion processes.
@@
The field of Generative AI has sparked significant excitement due to its broad range of applications, including both anticipated ones valued in the tens of billions of dollars and substantial academic and industrial research efforts (Bubeck et al., 2023; Abdin et al., 2023; Jiang et al., 2023; Stanford Institute for Human-Centered Artificial Intelligence, 2023). This expansive field encompasses thousands of models and tools that are driving research across various academic disciplines. However, alongside this enthusiasm, there is a pressing need for a more critical examination of the ethical implications associated with these tools and methods (Barnett, 2023). The deployment of these tools beyond academia has resulted in adverse consequences such as personal and financial losses, widespread information manipulation, and disruptions to sociotechnical infrastructure (Ferrara, 2024; Walker et al., 2023).
@@
For the purposes of our paper, we define Generative AI as being a subset of tools or processes that combine massive probability distributions with sense-making algorithms in order to create output that, taken as a whole, was not included in the training data (Jiang et al., 2023). The probability distribution should be based on the training data. The sense-making algorithm could be based on several strategies, such as a physical process like diffusion or an architectural approach like transformers/neural nets, although in the end, it should produce something akin to a latent/embedding space; a place to hold meaning (Gozalo-Brizuela and Garrido-Merchán, 2023; Zhao et al., 2023; Grand et al., 2022). Research is ongoing into the various distributions and processes that can be implemented in Generative AI models to achieve specific behaviors and results (Xu et al., 2023). However, irrespective of details, all of these models use advanced architectures in conjunction with large amounts of data gathered through brute-force methods across vast swathes of the publicly accessible internet (Xu et al., 2023; Zhao et al., 2023).
@@
In this paper, we are interested in how artists perceive Generative AI, what its impact on art-related careers might be, and what moral, ethical, artistic, and financial implications are coming hand-in-hand with the broader flourishing of Generative AI (Stanford Institute for Human-Centered Artificial Intelligence, 2023; Gozalo-Brizuela and Garrido-Merchán, 2023; Zhao et al., 2023). Therefore, we touch on three relevant topics – first, related academic work; second, the economic context for this discussion; and third, what sort of stages are within the pipeline of Generative AI models and how those stages can each cause harm (a specific and noteworthy type of impact) – within this Background section.
@@
Although in our paper we focus more on the pressing potential for harm, it is worth noting that there is also growing excitement around Generative AI’s capabilities and how it could be incorporated into artistic processes (Karson, 2023). For example, artist Cosmo Wenman worked with blind researchers Brandon Biggs, Lindsay Yazzolino, and Joshua Miele to create images using the Midjourney tool. Their team noted that programs like Midjourney can find “weird and thematic intersections” in art that go unnoticed by human eyes. They also discussed the role of chaos in creating “visual intrigue” and how technology can be used to “enhance agency.” Wenman called the new technology “essentially moveable type for visual rhetoric” (Schrader, 2023).
2.2. Related work
@@
A paper by Jiang et al. (Jiang et al., 2023) analyzes Generative AI art models and their potential impact on the world through the lens of diverse academic sources (both philosophical and scientific). They describe many opportunities for harm as a result of such products: violation of the consent and privacy of artists (Ho, 2024), a lack of dataset transparency (shielding violations and biases) (Ho, 2024), a failure to uphold the contextual integrity of content (Nissenbaum, 2004), harm to the livelihood of the artist, biases in the training data, automating processes in a way that is detrimental to the creation of art, reduced accessibility, erasure of human works by obscurity, and increasing reluctance by artists to share their works openly.
@@
The harms Jiang et al. (Jiang et al., 2023) foresee for artists fall into two categories, financial and emotional, which manifest as the replacement of a human artist on a project, the impersonation of an artist without consent (Gozalo-Brizuela and Garrido-Merchán, 2023), or plagiarism. Impersonation, in this context, could be weaponized for nefarious purposes (Ho, 2024). Moreover, these two kinds of harm, enacted against individuals at scale, create societal-level harm. Specific societal harms Jiang et al. discuss in their work include data laundering – through academic or open-source projects where public funds are funneled into research with private profits – the proliferation of stereotypes and “hegemonic views,” and erasure by obscurity (i.e., that AI-generated outputs can be created much more quickly than human ones). They also point out that US copyright and intellectual property jurisprudence needs to be better equipped to deal with these new tools and that open sourcing, often called on as a solution to software-related problems, is incomplete in this case (Jiang et al., 2023).
@@
As possible solutions, Jiang et al. bring up regulations regarding opt-in consent (for training data), independent funding (much current research is supported by industry), and education about how power interacts with technology. In particular, they note there are many incentives to label work as “apolitical” and that the spin capitalism is putting on Generative AI is the “democratization of art” (and other fields) (Jiang et al., 2023).
@@
Going beyond Jiang et al.’s list of solutions, Shan et al. (Shan et al., 2023) propose tools allowing artists to “poison” models that scrape their work without permission. The tool they created (Nightshade) can “destabilize general features” in the targeted generative model and “bleed-through” to semantically similar prompts without injecting large amounts of tailored data (Shan et al., 2023).
2.3. Economic context: livelihood of the artist
@@
AI’s integration into the global economy is rapidly expanding, as evidenced by the increasing number of AI-related job postings across various economic sectors (Stanford Institute for Human-Centered Artificial Intelligence, 2023; Squicciarini and Nachtigall, 2021). Employers worldwide are seeking AI-skilled workers, with many companies opting to adopt AI tools each year. The primary benefit reported by these companies is a reduction in costs, often achieved through layoffs and the substitution of human roles with AI-driven solutions (Stanford Institute for Human-Centered Artificial Intelligence, 2023; Gmyrek et al., 2023; Alekseeva et al., 2021).
@@
Some of these impacts are already coming to bear on the economic reality of artists (Chow, 2020; Henderson, 2010). Entry-level positions in film, TV, and gaming seem particularly at risk, and the conceptual period of a project seems to be a popular phase in which to replace artists with AI. Traditional niches like illustrating book covers can now be filled by the output of Generative AI, and major employers can reduce the number of employees needed (for example, Netflix Japan’s layoff of animators) (Jiang et al., 2023).
@@
The economic impact on artists is not solely determined by the output of AI models. Harm, both financial and emotional(Caporusso et al., 2023), can arise from the use of content without explicit consent. This includes situations where individuals may not have anticipated their work being repurposed by machine learning models or the entities controlling and utilizing these technologies (Nissenbaum, 2004). For instance, consider an artist with a physical gallery presence whose work is photographed by a visitor and subsequently used in ways they did not foresee, especially in the intricate landscape of social media (Ho, 2024).
@@
Moreover, the issue of copyright protection becomes paramount when AI models are trained using human artworks without proper consent, attribution, or adequate compensation for the original creators. This concern is shared among artists, creative industries, and government agencies alike (US Federal Trade Commission, 2023; Marcus and Southen, 2024).
2.4. Problems in the pipeline of Generative AI tools
@@
As discussed earlier, when assessing the impact of an AI model, it extends beyond the usage of the model’s output or its deployment (such as replacing human workers). Every stage involved in developing a Generative AI model can influence the eventual harm caused by the model, thereby affecting its overall impact. Although many stages in this pipeline are typically inaccessible, except for the input and output phases, they still play a significant role. In this section, we outline the pipeline of creating a Generative AI model and focus on one prevalent issue, bias (Zhao et al., 2023), to illustrate how various stages within the pipeline can contribute to the model’s impact.
@@
When developing an AI model, several steps are involved. Firstly, training materials need to be collected, often comprising a subset of real-world artifacts. These materials then undergo an assessment to address issues like poor quality, irrelevance, or over-representation of certain artifacts (Luccioni and Viviano, 2021; Zhao et al., 2023). Next, the model’s architecture is assembled, followed by the training process. For complex Generative AI models with unpredictable behavior, there is an additional step called alignment. This step aims to constrain the model’s range of potential behaviors to align with the creators’ desired outcomes (Zhao et al., 2023).
@@
Early in the pipeline, the selection of training material significantly shapes the eventual behavior of the model. For instance, many important datasets used in machine learning benchmarks predominantly feature lighter-skinned subjects (Buolamwini and Gebru, 2018), which indirectly introduces bias into the model. This type of global bias cannot be detected by locally examining each piece of training data for harmful content. Furthermore, datasets often contain objectionable and biased content, including racist hate speech and sexually inappropriate images (Luccioni and Viviano, 2021), even after undergoing cleaning procedures.
@@
In the middle of the pipeline, within the internal workings of the trained model, even fundamental and abstract elements like the meanings of common English words have been discovered to encode gender biases (Bolukbasi et al., 2016). As we progress toward the end of the pipeline, the output generated by the model can exhibit biases as well. For example, the Stable Diffusion model disproportionately generates images of white male executives, doctors, lawyers, judges, and law enforcement, and also disproportionately generates images of black and brown criminals and service workers (Ghosh and Caliskan, 2023). Additionally, users have reported that Generative AI tools often generate images where they appear younger, more sexualized, or with lighter skin tones, even when these attributes were not requested or suggested in the input prompt (Ghosh and Caliskan, 2023). These examples are not unique; similar biases are found across current AI models and products – including those of Generative AI – at every level, from the training data they learned from to their internal representations to their outputs (Luccioni and Viviano, 2021; Ghosh and Caliskan, 2023; Cheng et al., 2023; Zhao et al., 2023).
@@
As a final problem in the pipeline, it is worth noting that these AI tools cannot be created ex nihilo: they are integrated into complex human systems that bring their own forces to bear on the model’s potential impact. There are discrepancies in the demographics of who is wielding the power of these new technologies: who is making them (Stanford Institute for Human-Centered Artificial Intelligence, 2023), profiting off them, using them. This can have many knock-on effects, sometimes through mechanisms difficult to foresee (Even-Tov et al., 2023), with respect to the design, use, and role of these tools in society (Jiang et al., 2023; Stanford Institute for Human-Centered Artificial Intelligence, 2023). In particular, the companies behind these products often employ a great deal of secrecy as to their training data, architecture, or implementation (or all three) (Ho, 2024). As a result, users must interact with the product as a black box (Jiang et al., 2023; Marcus and Southen, 2024); concerned users can only conjecture as to what kind of training data could produce the observed results based on what is known of the basic architecture of the model, which in turn allows argument over the obscure details. The lack of transparency benefits the companies offering these products (with respect to establishing the impact, specifically the harms, of the model) in giving them a way to sidestep moral and ethical concerns (Marcus and Southen, 2024).
3. Methods
3.1. Survey Methodology

### v5

@@ Abstract
Introduction
Art has a rich history in co-evolution with the development of new technologies (Etro and Galenson 2023). The advent of Artificial Intelligence (AI) tools used to create or assist in the creative process carries great potential utility for artists but also harbors the significant potential for harm to the art workforce and artistic and intellectual property. To examine this tension between utility and threat, we surveyed 459 artists about their views on AI-generated art. The survey results reveal the nuanced perspectives of artists on Generative AI. When the discussion is thorough and directly related to their art practice, it prevents their views from being reduced to a simplistic binary sentiment, which fails to represent the diversity of their field. By capturing artists’ views through a survey, this method moves beyond anecdotal evidence and provides a counterbalance to third-party opinions as well as those of the Generative AI industry.
@@ Introduction
Background
There is ongoing disagreement regarding whether the output of Generative AI can be considered art (US Federal Trade Commission 2023; Jiang et al. 2023); we will not examine the topic in depth here. For convenience, we’ll refer to AI-generated output as “art” in this paper without asserting or refuting its artistic merit.
Jiang et al. define art as a human endeavor linked to culture and experience (Jiang et al. 2023). They argue that human creativity, distinguished by unique inspiration and corporeal engagement, differs from AI-generated work. Recent court decisions in the US have similarly denied copyright protection to autonomously produced AI creations, underscoring the role of human creativity in copyright law (US District Court for the District of Columbia 2023).
Three compelling arguments against considering AI output as art emerge: art’s human origin devoid of ”organic pressure,” the intimate connection between artist and world, and art’s unique role in human experience (Jiang et al. 2023; Huyghe 1970; Pfeifer and Bongard 2006). Regardless of the artistry debate, AI-generated creations can meet commercial standards, like book covers, entering diverse artistic spheres, and impacting practitioners.
Artistic use of technology and AI tools (Cetinic and She 2022) is not novel; artists historically adopt tools, including machines and automation, often innovating beyond standard applications (Huyghe 1970). Artists engaging with technology question its role in art and explore its boundaries, raising queries about intentionality, distraction, and artistic agency (Karson 2023).
It is important to note that, although how technology should be incorporated into art may be up for debate, that technology is part of the artistic process is not new or ambiguous (Mazzone and Elgammal 2019). Artists have always used tools (Huyghe 1970), and many art forms involve machines (such as woodwork, metalwork, and digital art) and automation (such as digital printers and cameras). Some artists also choose to use technology in ways other than its intended purpose or develop new art-making techniques by breaking technology. Such artists may be both suspicious and intrigued by the technology they use, questioning the relationship between humans and machines and expanding art practices and processes. When presented with new technology, an artist may ask questions like “How can I keep my work with the machine intentional? How might this technology distract me? Is the machine using me, or am I using it? Is this image-making or art-making? How does this automation serve my artistic intentions and practice?” (Karson 2023).
@@
Generative AI models can generate outputs across various media types based on user inputs, such as text prompts or multimedia (Gozalo-Brizuela and Garrido-Merchán 2023). Key players in textual Generative AI include large language models (LLMs) like OpenAI’s ChatGPT and GPT-4, while visual models like StableDiffusion and Midjourney modify images using iterative noise adjustments guided by diffusion processes.
@@
The field of Generative AI has sparked significant excitement due to its broad range of applications, including both anticipated ones valued in the tens of billions of dollars and substantial academic and industrial research efforts (Bubeck et al. 2023; Abdin et al. 2023; Jiang et al. 2023; Stanford Institute for Human-Centered Artificial Intelligence 2023). This expansive field encompasses thousands of models and tools that are driving research across various academic disciplines. However, alongside this enthusiasm, there is a pressing need for a more critical examination of the ethical implications associated with these tools and methods (Barnett 2023). The deployment of these tools beyond academia has resulted in adverse consequences such as personal and financial losses, widespread information manipulation, and disruptions to sociotechnical infrastructure (Ferrara 2024; Walker et al. 2023).
@@
For the purposes of our paper, we define Generative AI as being a subset of tools or processes that combine massive probability distributions with sense-making algorithms in order to create output that, taken as a whole, was not included in the training data (Jiang et al. 2023). The probability distribution should be based on the training data. The sense-making algorithm could be based on several strategies, such as a physical process like diffusion or an architectural approach like transformers/neural nets, although in the end, it should produce something akin to a latent/embedding space; a place to hold meaning (Gozalo-Brizuela and Garrido-Merchán 2023; Zhao et al. 2023; Grand et al. 2022). Research is ongoing into the various distributions and processes that can be implemented in Generative AI models to achieve specific behaviors and results (Xu et al. 2023). However, irrespective of details, all of these models use advanced architectures in conjunction with large amounts of data gathered through brute-force methods across vast swathes of the publicly accessible internet (Xu et al. 2023; Zhao et al. 2023).
@@
In this paper, we are interested in how artists perceive Generative AI, what its impact on art-related careers might be, and what moral, ethical, artistic, and financial implications are coming hand-in-hand with the broader flourishing of Generative AI (Stanford Institute for Human-Centered Artificial Intelligence 2023; Gozalo-Brizuela and Garrido-Merchán 2023; Zhao et al. 2023). Therefore, we touch on three relevant topics – first, related academic work; second, the economic context for this discussion; and third, what sort of stages are within the pipeline of Generative AI models and how those stages can each cause harm (a specific and noteworthy type of impact) – within this Background section.
@@
Although in our paper we focus more on the pressing potential for harm, it is worth noting that there is also growing excitement around Generative AI’s capabilities and how it could be incorporated into artistic processes (Karson 2023). For example, artist Cosmo Wenman worked with blind researchers Brandon Biggs, Lindsay Yazzolino, and Joshua Miele to create images using the Midjourney tool. Their team noted that programs like Midjourney can find “weird and thematic intersections” in art that go unnoticed by human eyes. They also discussed the role of chaos in creating “visual intrigue” and how technology can be used to “enhance agency.” Wenman called the new technology “essentially moveable type for visual rhetoric” (Schrader 2023).
Related Work
@@
A paper by Jiang et al. (Jiang et al. 2023) analyzes Generative AI art models and their potential impact on the world through the lens of diverse academic sources (both philosophical and scientific). They describe many opportunities for harm as a result of such products: violation of the consent and privacy of artists (Ho 2024), a lack of dataset transparency (shielding violations and biases) (Ho 2024), a failure to uphold the contextual integrity of content (Nissenbaum 2004), harm to the livelihood of the artist, biases in the training data, automating processes in a way that is detrimental to the creation of art, reduced accessibility, erasure of human works by obscurity, and increasing reluctance by artists to share their works openly.
@@
The harms Jiang et al. (Jiang et al. 2023) foresee for artists fall into two categories, financial and emotional, which manifest as the replacement of a human artist on a project, the impersonation of an artist without consent (Gozalo-Brizuela and Garrido-Merchán 2023), or plagiarism. Impersonation, in this context, could be weaponized for nefarious purposes (Ho 2024). Moreover, these two kinds of harm, enacted against individuals at scale, create societal-level harm. Specific societal harms Jiang et al. discuss in their work include data laundering – through academic or open-source projects where public funds are funneled into research with private profits – the proliferation of stereotypes and “hegemonic views,” and erasure by obscurity (i.e., that AI-generated outputs can be created much more quickly than human ones). They also point out that US copyright and intellectual property jurisprudence needs to be better equipped to deal with these new tools and that open sourcing, often called on as a solution to software-related problems, is incomplete in this case (Jiang et al. 2023).
@@
As possible solutions, Jiang et al. bring up regulations regarding opt-in consent (for training data), independent funding (much current research is supported by industry), and education about how power interacts with technology. In particular, they note there are many incentives to label work as “apolitical” and that the spin capitalism is putting on Generative AI is the “democratization of art” (and other fields) (Jiang et al. 2023).
@@
Going beyond Jiang et al.’s list of solutions, Shan et al. (Shan et al. 2023) propose tools allowing artists to “poison” models that scrape their work without permission. The tool they created (Nightshade) can “destabilize general features” in the targeted generative model and “bleed-through” to semantically similar prompts without injecting large amounts of tailored data (Shan et al. 2023).
Economic Context: Livelihood of the Artist
@@
AI’s integration into the global economy is rapidly expanding, as evidenced by the increasing number of AI-related job postings across various economic sectors (Stanford Institute for Human-Centered Artificial Intelligence 2023; Squicciarini and Nachtigall 2021). Employers worldwide are seeking AI-skilled workers, with many companies opting to adopt AI tools each year. The primary benefit reported by these companies is a reduction in costs, often achieved through layoffs and the substitution of human roles with AI-driven solutions (Stanford Institute for Human-Centered Artificial Intelligence 2023; Gmyrek, Berg, and Bescond 2023; Alekseeva et al. 2021).
@@
Some of these impacts are already coming to bear on the economic reality of artists (Chow 2020; Henderson 2010). Entry-level positions in film, TV, and gaming seem particularly at risk, and the conceptual period of a project seems to be a popular phase in which to replace artists with AI. Traditional niches like illustrating book covers can now be filled by the output of Generative AI, and major employers can reduce the number of employees needed (for example, Netflix Japan’s layoff of animators) (Jiang et al. 2023).
@@
The economic impact on artists is not solely determined by the output of AI models. Harm, both financial and emotional(Caporusso et al. 2023), can arise from the use of content without explicit consent. This includes situations where individuals may not have anticipated their work being repurposed by machine learning models or the entities controlling and utilizing these technologies (Nissenbaum 2004). For instance, consider an artist with a physical gallery presence whose work is photographed by a visitor and subsequently used in ways they did not foresee, especially in the intricate landscape of social media (Ho 2024).
@@
Moreover, the issue of copyright protection becomes paramount when AI models are trained using human artworks without proper consent, attribution, or adequate compensation for the original creators. This concern is shared among artists, creative industries, and government agencies alike (US Federal Trade Commission 2023; Marcus and Southen 2024).
Problems in the Pipeline of Generative AI Tools
@@
As discussed earlier, when assessing the impact of an AI model, it extends beyond the usage of the model’s output or its deployment (such as replacing human workers). Every stage involved in developing a Generative AI model can influence the eventual harm caused by the model, thereby affecting its overall impact. Although many stages in this pipeline are typically inaccessible, except for the input and output phases, they still play a significant role. In this section, we outline the pipeline of creating a Generative AI model and focus on one prevalent issue, bias (Zhao et al. 2023), to illustrate how various stages within the pipeline can contribute to the model’s impact.
@@
When developing an AI model, several steps are involved. Firstly, training materials need to be collected, often comprising a subset of real-world artifacts. These materials then undergo an assessment to address issues like poor quality, irrelevance, or over-representation of certain artifacts (Luccioni and Viviano 2021; Zhao et al. 2023). Next, the model’s architecture is assembled, followed by the training process. For complex Generative AI models with unpredictable behavior, there is an additional step called alignment. This step aims to constrain the model’s range of potential behaviors to align with the creators’ desired outcomes (Zhao et al. 2023).
@@
Early in the pipeline, the selection of training material significantly shapes the eventual behavior of the model. For instance, many important datasets used in machine learning benchmarks predominantly feature lighter-skinned subjects (Buolamwini and Gebru 2018), which indirectly introduces bias into the model. This type of global bias cannot be detected by locally examining each piece of training data for harmful content. Furthermore, datasets often contain objectionable and biased content, including racist hate speech and sexually inappropriate images (Luccioni and Viviano 2021), even after undergoing cleaning procedures.
@@
In the middle of the pipeline, within the internal workings of the trained model, even fundamental and abstract elements like the meanings of common English words have been discovered to encode gender biases (Bolukbasi et al. 2016). As we progress toward the end of the pipeline, the output generated by the model can exhibit biases as well. For example, the Stable Diffusion model disproportionately generates images of white male executives, doctors, lawyers, judges, and law enforcement, and also disproportionately generates images of black and brown criminals and service workers (Ghosh and Caliskan 2023). Additionally, users have reported that Generative AI tools often generate images where they appear younger, more sexualized, or with lighter skin tones, even when these attributes were not requested or suggested in the input prompt (Ghosh and Caliskan 2023). These examples are not unique; similar biases are found across current AI models and products – including those of Generative AI – at every level, from the training data they learned from to their internal representations to their outputs (Luccioni and Viviano 2021; Ghosh and Caliskan 2023; Cheng, Durmus, and Jurafsky 2023; Zhao et al. 2023).
@@
As a final problem in the pipeline, it is worth noting that these AI tools cannot be created ex nihilo: they are integrated into complex human systems that bring their own forces to bear on the model’s potential impact. There are discrepancies in the demographics of who is wielding the power of these new technologies: who is making them (Stanford Institute for Human-Centered Artificial Intelligence 2023), profiting off them, using them. This can have many knock-on effects, sometimes through mechanisms difficult to foresee (Even-Tov et al. 2023), with respect to the design, use, and role of these tools in society (Jiang et al. 2023; Stanford Institute for Human-Centered Artificial Intelligence 2023). In particular, the companies behind these products often employ a great deal of secrecy as to their training data, architecture, or implementation (or all three) (Ho 2024). As a result, users must interact with the product as a black box (Jiang et al. 2023; Marcus and Southen 2024); concerned users can only conjecture as to what kind of training data could produce the observed results based on what is known of the basic architecture of the model, which in turn allows argument over the obscure details. The lack of transparency benefits the companies offering these products (with respect to establishing the impact, specifically the harms, of the model) in giving them a way to sidestep moral and ethical concerns (Marcus and Southen 2024).
Methods
Survey Methodology